OpenStudy (anonymous):

A 500-g sample of aluminum sulfate is reacted with 450 g of calcium hydroxide. If 500 grams of calcium sulfate are produced, what is the percent yield of the reaction? Please show all work

OpenStudy (abmon98):

Aluminium has a charge of +3 and So4-2 ion interchange the charge to form Al2(SO4)^3 Calcium +2 and OH- interchane Charges Ca(OH)2 Al2(SO4)3+Ca(OH)2-->CaSO4+Al(OH)3 Balance :Write down the number of atoms that you have on each side of the equation. On the Reactant Side: Al:2 S:3 O:14 H:2 On the Product Side: Al:1 S:1 O:7 H:3 Add a coefficient of 2 Al(OH)3 Al2(SO4)3+Ca(OH)2-->CaSO4+2Al(OH)3 On the Reactant Side: Al:2 S:3 O:14 H:2 On the Product Side: Al:2 S:1 O:10 H:6 Add a Coefficient of 3 to CaSO4 Al2(SO4)3+Ca(OH)2-->3CaSO4+2Al(OH)3 On the Reactant Side: Al:2 S:3 O:14 H:2 Ca:1 On the Product Side: Al:2 S:3 O:18 H:6 Ca:3 Add a coeffcient of 3 to Ca(OH)2 Al2(SO4)3+3Ca(OH)2-->3CaSO4+2Al(OH)3 On the Reactant Side: Al:2 S:3 O:18 H:6 Ca:3 On the Product Side: Al:2 S:3 O:18 H:6 Ca:3 Write out each mass given if 500 grams of CaSO4 are produced this is the actual yield 500*1mole of Al2(SO4)3/342.15 Al2(SO4)3=1.46 Moles*3 moles of CaSO4/1mole of Al2(SO4)3=4.38*136=595.68 if 450 *1 mole/74.=6.08 mole*3 moles of CaSO4/3 Moles of Ca(OH)2=6.08*136=827.02 grams Choose the smaller Number (Actual Yield/Theortical Yield)*100 (500/595.68)*100=84.93

OpenStudy (anonymous):

wow, thank you!

OpenStudy (abmon98):

your welcome :)